OpenStudy (anonymous):

What is the best order-of-magnitude estimate for 789,000? A. 10(7) B. 10(6) C. 10(5) D. 10(4)

OpenStudy (anonymous):

i think b @31356

OpenStudy (anonymous):

10 * 10 = 100 100 * 10 = 1000 1000 * 10 = 10000 1000 * 10 = 100000 100000 * 10 = 1000000 1000000 is closest to 789,000?

OpenStudy (31356):

You are correct!

OpenStudy (anonymous):

can u just explain how? i was not sure if it was b or c

OpenStudy (31356):

Ok..

OpenStudy (anonymous):

i said b because i rounded, was that necessrasy?

OpenStudy (anonymous):

nvm i got it i have one more mind helping???

OpenStudy (anonymous):

here is medal btw

OpenStudy (31356):

Oh ok I was going to explain it but ok.

OpenStudy (31356):

Sure..

OpenStudy (31356):

Thgx for the medal ;)